alright, so I am doing a pie chart for grade 9 geography and I don't know how to get my data into percentages for the pie chart. my data numbers are,
-544
-110
-81
how do I fit these into a pie chart

Answer by checkley77(12844)

-544
-110
-81
----------
735 is the total value of the pie.
544/735=74% of the pie.
110/735=15% of the pie.
81/735=11% of the pie.
74+15+11=100% of the pie.